OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(-5, 1), C(0, 6), D(5, 1) and E (0, -4) by showing that all four sides are equal. A. BD = CE = square root of 50 B. BD = CE = 50 C. BC = CD = DE = EB = 50 D. BC = CD = DE = EB = square root of 50

OpenStudy (anonymous):

D. B(-5, 1), C(0, 6) \[ \sqrt{ (6-1)^2+(0-(-5))^2 }=\sqrt{50} \]

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

YW